OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. y - 2 = 4(x + 6) y + 6 = 4(x - 2) y + 2 = 4(x - 6) y - 6 = 4(x + 2)

OpenStudy (anonymous):

y-y1=m(x-x1) m=slope

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

the last one